### v1

# Qu, Galichon, Ugander

[…]
Qu, Galichon, Ugander
[…]
Sinkhorn’s Algorithm and Choice Modeling
[…]
On Sinkhorn’s Algorithm and Choice Modeling
[…]
1 Introduction
The modeling of choice and ranking data is an important topic across many disciplines. Given a collection of $m$ objects, a universal problem is to aggregate choice or partial ranking data over them to arrive at a reasonable description of either the behavior of decision makers, the intrinsic qualities of the objects, or both. Work on such problems dates back over a century at least to the work of Landau, who considered $m$ chess players and a record of their match results against one another, aiming to aggregate the pairwise comparisons to arrive at a global ranking of all players (Landau 1895, Elo 1978). More generally, comparison data can result from choices from subsets of varying sizes, from partial or complete rankings of objects, or from mixtures of different data types.
The modern rigorous study of comparisons primarily builds on the foundational works of Thurstone (1927) and Zermelo (1929). Both proposed models based on a numerical “score” for each item (e.g., chess player), but with different specifications of choice probabilities. Zermelo (1929) builds on the intuition that choice probability should be proportional to the score, and proposes a iterative algorithm to estimate the scores from pairwise comparison data. As one of the foundational works in this direction, Luce (1959) formalized the multinomial logit model of discrete choice starting from the axiom of independence of irrelevant alternatives (IIA). It states that the relative likelihood of choosing an item $j$ over another item $k$ is independent of the presence of other alternatives. In other words, if $S$ and $S^{\prime}$ are two subsets of the $m$ alternatives, both containing $j$ and $k$, and $\Pr(j,S)$ denotes the probability of choosing item $j$ from $S$, then for $\Pr(k,S)>0$,
$\frac{\Pr(j,S)}{\Pr(k,S)}$ $=\frac{\Pr(j,S^{\prime})}{\Pr(k,S^{\prime})}.$
This invariance property, together with a condition for zero probability alternatives, are often referred to as Luce’s choice axioms. They guarantee that each alternative can be summarized by a non-negative score $s_{j}$ such that the probability of choice can be parameterized by
$\Pr(j,S)$ $=\frac{s_{j}}{\sum_{k\in S}s_{k}},$ (1)
for any set $S$ that contains $j$. The parameters $s_{j}$ reflect the “intrinsic” value of item $i$, and are unique up to a normalization, which can be set to $\sum_{j}s_{j}=1$. This choice model generalizes the Bradley–Terry–Luce model (BTL) for pairwise comparisons (Bradley and Terry 1952), and also applies to ranking data when each $k$-way ranking is broken down into $k-1$ choice observations, where an item is chosen over the set of items ranked lower (Critchlow et al. 1991, Plackett 1975, Hausman and Ruud 1987). The many subsequent works that build on Luce’s choice axioms speak to its fundamental importance in choice modeling. Other works have also sought to address the limitations of Luce choice models. Prominent among them are probit models (Thurstone 1927, Berkson 1944), random utility (RUM) models (McFadden and Train 2000), context-dependent (CDM) models (Batsell and Polking 1985, Seshadri et al. 2020), and behavioral models from psychology (Tversky 1972).
Matrix balancing, meanwhile, is a seemingly unrelated mathematical problem. In its most common form (what we study in this paper), the problem seeks positive column and row scalings $D^{0},D^{1}$ of a non-negative matrix $A$, such that the scaled matrix $D^{0}AD^{1}$ has row and column sums equal to some prescribed positive marginals $p,q$ (Sinkhorn 1967). Over the years, numerous applications and problems across different domains, including statistics (Yule 1912, Deming and Stephan 1940), economics (Stone 1962, Galichon 2021), transportation and networks (Kruithof 1937, Lamond and Stewart 1981, Chang et al. 2021), optimization (Ruiz 2001), and machine learning (Cuturi 2013, Peyré et al. 2019), have found themselves essentially solving a new incarnation of the old matrix balancing problem, which attests to its universality and importance.
A major appeal of the matrix balancing problem lies in the simplicity and elegance of its popular solution method, Sinkhorn’s algorithm. It simply alternates between updating the scalings to satisfy one of the two marginal conditions, leading to lightweight implementations that have proved extremely effective for practical problems. Despite this widespread use, the convergence behavior of Sinkhorn’s algorithm has still not been fully understood. In particular, while there have been extensive studies of convergence, many assume that the matrix $A$ is (entry-wise) strictly positive. In contrast, quantitative analyses for the case when $A$ contains zeros entries are rare and fragmented, employing different assumptions whose connections and distinctions remain unclear. For example, the recent work of Léger (2021) establishes a global sub-linear $\mathcal{O}(1/t)$ error bound, where $t$ is the number of iterations, under minimal assumptions that guarantee convergence. On the other hand, for square matrix $A$ and uniform marginals $p,q$, Knight (2008) shows that the convergence is asymptotically linear whenever finite scalings $D^{0},D^{1}$ exist, i.e., errors $e^{(t)}$ satisfy $e^{(t+1)}/e^{(t)}\rightarrow\lambda$ for some $\lambda\in(0,1)$. It has remained as important open problems to establish quantitative global geometric (linear) convergence with an $\mathcal{O}(\lambda^{t})$ error bound, as well as sharp asymptotic bounds for Sinkhorn’s algorithm that are applicable to general non-negative $A$ and non-uniform $p,q$. Moreover, it remains to clarify the convergence behavior under different assumptions on the problem structure.
In this paper, we resolve several long-standing convergence questions. Surprisingly, the inspirations for our results come from connections to choice modeling and seemingly independent results in that domain. Our first contribution is in recognizing choice modeling a la Luce as yet another instance where a central problem reduces to that of matrix balancing. We clarify the equivalence between problem assumptions in each setting. Furthermore, we demonstrate that classic and new algorithms from the choice literature, including those of Zermelo (1929), Dykstra (1956), Ford (1957), Berry et al. (1995), Hunter (2004), Maystre and Grossglauser (2017), Agarwal et al. (2018), can be viewed as variants of or analogous to Sinkhorn’s algorithm.
These intimate mathematical and algorithmic connections allow us to provide a unifying perspective on works from both areas. More importantly, they enable researchers to import insights and tools from one domain to the other. In particular, recent works on choice modeling (Shah et al. 2015, Seshadri et al. 2020, Vojnovic et al. 2020) have highlighted the importance of algebraic connectivity of data for estimation and computation, which motivates us to consider this measure in the analysis of Sinkhorn’s algorithm.
Our next contribution is to obtain a global $\mathcal{O}(\lambda^{t})$ linear convergence bound for Sinkhorn’s algorithm, whenever the matrix balancing problem has a finite solution pair $D^{0},D^{1}$. Moreover, the convergence rate $\lambda$ depends on the algebraic connectivity of a bipartite graph associated with $A$. This result complements the linear convergence result of Franklin and Lorenz (1989) for positive matrices, and the sub-linear bound of Léger (2021) for non-negative matrices under weaker conditions. In addition, we also characterize the asymptotic linear rate of convergence in terms of the solution matrix $D^{0}AD^{1}$ and target marginals $p,q$, directly generalizing the result of Knight (2008), but with a more explicit analysis that exploits an intrinsic orthogonality structure of Sinkhorn’s algorithm. The convergence analyses and results developed in this paper apply irrespective of whether $A$ is strictly positive or not, making them applicable to a wide array of settings. They also clarify the convergence behavior of Sinkhorn’s algorithm under two regimes: when a finite scaling exists, Sinkhorn’s algorithm always converges linearly; otherwise, it only converges sub-linearly under the minimum conditions required for convergence.
Our convergence results highlight the importance of algebraic connectivity and related spectral properties for the convergence of Sinkhorn’s algorithm. At first glance, this dependence may seem unintuitive. However, once we interpret Sinkhorn’s algorithm as a distributed optimization algorithm on a bipartite graph (Section 9), it is less surprising that the convergence behavior is governed by the spectral properties of that graph.
The challenges of ill-defined matrix balancing problems and non-convergence of Sinkhorn’s algorithm in practice also motivate us to propose a regularized version of Sinkhorn’s algorithm. It is inspired by the regularization of Luce choice models using Gamma priors, and is guaranteed to converge even when the necessary condition for the convergence of the standard algorithm does not hold. This regularized Sinkhorn’s algorithm could be useful in practice, especially in applications with very sparse data.
We believe that the connections we establish in this paper between choice modeling and matrix balancing can lead to more interesting results in both directions, and are therefore relevant to researchers from both communities.
2 Related Work
This section includes an extensive review of related works in choice modeling and matrix balancing. Well-versed readers may skip ahead to the mathematical preliminaries (Section 3) and our core results (Sections 4 and 5).
2.1 Choice Modeling
Methods for aggregating choice and comparison data usually take one of two closely related approaches: maximum likelihood estimation of a statistical model or ranking according to the stationary distributions of a random walk on a Markov chain. Recent connections between maximum likelihood and spectral methods have put these two classes of approaches in increasingly close conversation with each other.
Spectral Methods. The most well-known spectral method for rank aggregation is perhaps the PageRank algorithm (Page et al. 1999), which ranks web pages based on the stationary distribution of a random walk on a hyperlink graph. The use of stationary distributions also features in the work of Dwork et al. (2001), the Rank Centrality (RC) algorithm (Negahban et al. 2012, 2016), which generates consistent estimates for the Bradley–Terry–Luce pairwise comparison model under assumptions on the sampling frame, and the Luce Spectral Ranking (LSR) and iterative LSR (I-LSR) algorithms of Maystre and Grossglauser (2015) for choices from pairs as well as larger sets. Following that work, Agarwal et al. (2018) proposed the Accelerated Spectral Ranking (ASR) algorithm with provably faster mixing times than RC and LSR, and better sample complexity bounds than Negahban et al. (2016). Knight (2008) is an intriguing work partially motivated by Page et al. (1999) that applies Sinkhorn’s algorithm, which is central to the current work, to compute authority and hub scores similar to those proposed by Kleinberg (1999) and Tomlin (2003), although the focus in Knight (2008) is on Markov chains rather than maximum likelihood estimation of choice models. For ranking data, Soufiani et al. (2013) decompose rankings into pairwise comparisons and develop consistent estimators for Plackett–Luce models based on a generalized method of moments. Other notable works that make connections between Markov chains and choice modeling include Blanchet et al. (2016) and Ragain and Ugander (2016).
Maximum Likelihood Methods. Maximum likelihood estimation of the Bradley–Terry–Luce model dates back to Zermelo (1929), Dykstra (1956), and Ford (1957), which all give variants of the same iterative algorithm and prove its convergence to the MLE when the directed comparison graph is strongly connected. Much later, Hunter (2004) observed that their algorithms are instances of the class of minorization-maximization (MM) algorithms, and develops MM algorithms for the Plackett–Luce model for ranking data, among others. Vojnovic et al. (2020) further investigate the convergence of the MM algorithm for choice models. Newman (2023) proposes an alternative to the classical iterative algorithm for pairwise comparisons based on a reformulated moment condition, achieving impressive empirical speedups. Negahban et al. (2012) is arguably the first work that connects maximum likelihood estimation to Markov chains, followed by Maystre and Grossglauser (2015), whose spectral method is based on a balance equation interpretation of the optimality condition. Kumar et al. (2015) consider the problem of inverting the stationary distribution of a Markov chain, and embed the maximum likelihood problem of the Luce choice model into this framework, where the MLEs parameterize the desired transition matrix. Maystre and Grossglauser (2017) consider the estimation of a network choice model with similarly parameterized random walks. Lastly, a vast literature in econometrics on discrete choice also considers different aspects of the ML estimation problem. In particular, the present paper is related to the Berry–Levinsohn–Pakes (BLP) framework of Berry et al. (1995), well-known in econometrics. The matrix balancing interpretation of maximum likelihood estimation of choice models that we develop in this paper connects many of the aforementioned works.
Besides the optimization of a given maximum likelihood problem, there have also been extensive studies of the statistical properties of the maximum likelihood estimator for these choice problems. Hajek et al. (2014) prove that the MLE is minimax optimal for $k$-way rankings, and Rajkumar and Agarwal (2014) show that the MLE for Bradley–Terry–Luce can recover the correct ranking under model mis-specification when noise is “bounded”. As a byproduct of analysis for a context-dependent generalization of the Luce choice model, Seshadri et al. (2020) obtain tight expected risk and tail risk bounds for the MLEs of Luce choice models (which they call MNL) and Plackett–Luce ranking models, extending and improving upon previous works by Hajek et al. (2014), Shah et al. (2015), Vojnovic and Yun (2016).
Our present work is primarily concerned with the optimization aspects of the maximum likelihood problem. Nonetheless, the importance of algebraic connectivity—as quantified by the Fiedler eigenvalue (Fiedler 1973)—in the results of Shah et al. (2015), Seshadri et al. (2020) as well as Vojnovic et al. (2020) provides motivations in our convergence analysis of Sinkhorn’s algorithm for matrix balancing.
Lastly, a short note on terminology. Even though a choice model based on (1) is technically a “multinomial logit model” with only intercept terms (McFadden et al. 1973), there are subtle differences. When (1) is applied to model ranking and choice data with distinct items, each observation $i$ usually consists of a possibly different subset $S_{i}$ of the universe of all alternatives, so that there is a large number of different configurations of the choice menu in the dataset. On the other hand, common applications of multinomial logit models, such as classification models in statistics and machine learning (Bishop and Nasrabadi 2006) and discrete choice models in econometrics (McFadden et al. 1973), often deal with repeated observations consisting of the same number of alternatives. However, these alternatives now possess “characteristics” that vary across observations, which are often mapped parametrically to the scores in (1). In this paper, we primarily use the term Luce choice model to refer to the model (3), although it is also called MNL (for multinomial logit) models in some works. We refrain from using the term MNL to avoid confusion with parametric models for featurized items used in ML and econometrics.
2.2 Matrix Balancing
The matrix balancing problem we study in this paper and its variants (Ruiz 2001, Bradley 2010) underlie a diverse range of applications from different disciplines. The question of scaling rows and columns of a matrix $A$ so that the resulting matrix has target row and column norms $p,q$ has been studied as early as the 1930s, and continue to intrigue researchers from different backgrounds. The present paper only contains a partial survey of the vast literature on the topic. Schneider and Zenios (1990) and Idel (2016) provide excellent discussions of many applications. In Section 11, we present concise summaries of some applications to illustrate the ubiquity of the matrix balancing problem.
The standard iterative algorithm for the matrix balancing problem has been rediscovered independently quite a few times. As a result, it has domain-dependent names, including the iterative proportional fitting (IPF) procedure (Deming and Stephan 1940), biproportional fitting (Bacharach 1965) and the RAS algorithm (Stone 1962), but is perhaps most widely known as Sinkhorn’s algorithm (Sinkhorn 1964, Cuturi 2013). A precise description can be found in Algorithm 1. This algorithm is also closely related to relaxation and coordinate descent type methods for solving the dual of entropy minimization problems (Bregman 1967b, Cottle et al. 1986, Tseng and Bertsekas 1987, Luo and Tseng 1992), as well as message passing and belief propagation algorithms in distributed optimization (Balakrishnan et al. 2004, Agarwal et al. 2018).
The convergence behavior of Sinkhorn’s algorithm has been extensively studied by Sinkhorn (1964), Bregman (1967a), Lamond and Stewart (1981), Franklin and Lorenz (1989), Ruschendorf (1995), Kalantari et al. (2008), Knight (2008), Pukelsheim and Simeone (2009), Altschuler et al. (2017), Chakrabarty and Khanna (2021), Léger (2021), among many others. For $A$ with strictly positive entries, Franklin and Lorenz (1989) first established the global linear convergence of Sinkhorn’s algorithm in the Hilbert projective metric $d$. In particular, if $r^{(t)}$ denotes the row sum of the scaled matrix after $t$ iterations of Sinkhorn’s algorithm that enforce column constraints, then
$d(r^{(t)},p)\leq\lambda^{t}\cdot d(r^{(0)},p)$ (2)
for some $\lambda\in(0,1)$ dependent on $A$. See Bushell (1973) for details on the Hilbert metric. Altschuler et al. (2017) and Chakrabarty and Khanna (2021) show global sub-linear convergence in terms of iteration complexity bounds on the $\ell^{1}$ distance independent of matrix dimension.
However, the matter of convergence is more delicate when the matrix contains zero entries, and additional assumptions on the problem structure are required. For non-negative $A$, convergence is first established by Sinkhorn and Knopp (1967) in the special case of square $A$ and uniform $p=q=\mathbf{1}_{n}=\mathbf{1}_{m}$. Their necessary and sufficient condition is that $A$ has total support, i.e., any non-zero entry of $A$ must be in $(A_{1\sigma(1)},A_{2\sigma(2)},\dots,A_{n\sigma(n)})$ for some permutation $\sigma$. Soules (1991) shows the convergence is linear, and Knight (2008) provides an explicit and tight asymptotic linear convergence rate in terms of the sub-dominant (second largest) singular value of the scaled doubly stochastic matrix $D^{0}AD^{1}$. However, no asymptotic linear convergence rate is previously known for non-uniform marginals.
For general non-negative matrices and non-uniform marginals, the necessary and sufficient conditions for the matrix balancing problem that generalize that of Sinkhorn and Knopp (1967) have been studied by Thionet (1964), Bacharach (1965), Brualdi (1968), Menon (1968), Djoković (1970), Sinkhorn (1974), Balakrishnan et al. (2004), Pukelsheim and Simeone (2009), and convergence of Sinkhorn’s algorithm under these conditions is well-known. Connecting Sinkhorn’s algorithm to dual coordinate descent for entropy minimization, Luo and Tseng (1992) show that the dual objective converges linearly with some unknown rate $\lambda$ when finite scalings $D^{0},D^{1}$ exist. However, their result is implicit and there are no results that quantify this rate $\lambda$, even for special classes of non-negative matrices. When convergence results on positive matrices in previous works are applied to non-negative matrices, the bounds often blow up or become degenerate as soon as $\min_{ij}A_{ij}\downarrow 0$. For example, in (2) the contraction factor $\lambda\rightarrow 1$ when $A$ contains zero entries. In contrast, under a weaker condition that guarantees the convergence of Sinkhorn’s algorithm, Léger (2021) gives a quantitative global $\mathcal{O}(1/t)$ bound for non-negative matrices. It remains to reconcile the results of these works and characterize the linear rate $\lambda$ for non-negative $A$.
Our work precisely fills the gaps left by these works. The global linear convergence result in Theorem 5.1 establishes a contraction like (2) whenever finite scalings $D^{0},D^{1}$ exist, and characterize $\lambda$ in terms of the algebraic connectivity. Moreover, the asymptotic linear rate in Theorem 5.4 directly extends the result of Knight (2008). See Table 1 for a detailed summary and comparison of the convergence results in previous works and this paper.
The dependence of Sinkhorn’s convergence rate on spectral properties of graphs can be compared to convergence results in the literature on decentralized optimization and gossip algorithms, where a spectral gap quantifies the convergence rate (Boyd et al. 2006, Xiao et al. 2007).
3 Preliminaries on Choice Modeling and Matrix Balancing
We start by providing brief but self-contained introductions to the two main subjects of this paper, choice modeling and matrix balancing, including their respective underlying mathematical problems and assumptions. Then we formally establish their equivalence in Section 4.
3.1 Maximum Likelihood Estimation of Luce Choice Models
In the Luce choice modeling framework, we have $n$ observations $\{(j_{i},S_{i})\}_{i=1}^{n}$, each consisting of a choice set $S_{i}\subseteq\{1,\dots,m\}=[m]$ that is a subset of the total $m$ alternatives/items/objects, and the alternative selected, denoted by $j_{i}\in S_{i}$. The choice probability is prescribed by Luce’s axiom of choice given model parameter $s\in\mathbb{R}^{m}_{++}$ in the interior of the probability simplex $\Delta_{m}$:
$\Pr(j_{i},S_{i})$ $=\frac{s_{j_{i}}}{\sum_{k\in S_{i}}s_{k}},$
and the likelihood of the observed data is thus given by
$L(s;\{(j_{i},S_{i})\}_{i=1}^{n}):=$ $\prod_{i=1}^{n}\frac{s_{j_{i}}}{\sum_{k\in S_{i}}s_{k}}.$ (3)
Importantly, the fact that (3) has potentially different choice sets $S_{i}$ across observations have structure-dependent implications for both the optimization and the statistical efficiencies of the model.
The log-likelihood of the choice model described by (3) is
$\ell(s):=\log L(s)$ $=\sum_{i=1}^{n}\log s_{j_{i}}-\log\sum_{k\in S_{i}}s_{k}.$ (4)
To estimate $s=\{s_{1},\dots,s_{m}\}$, we focus on the maximum likelihood estimation approach, which maximizes (4) over the interior of the probability simplex. If we reparameterize $\exp(u_{j})=s_{j}$, it is obvious that (4) is concave in $u$. However, to ensure the log-likelihood (4) has a unique maximizer in the interior of the simplex, additional assumptions on the comparison structure of the dataset $\{(j_{i},S_{i})\}_{i=1}^{n}$ are needed. The following classic condition is necessary and sufficient for the maximum likelihood problem to be well-posed. {assumption}[Strong Connectivity] In any partition of $[m]$ into two nonempty subsets $S$ and its complement $S^{C}$, some $j\in S$ is selected at least once over some $k\in S^{C}$. Equivalently, the directed comparison graph, with items as vertices and an edge $j\rightarrow k$ iff $k$ is selected in some $S_{i}$ for which $j,k\in S_{i}$, is strongly connected. Section 3.1 is standard in the literature (Hunter 2004, Noothigattu et al. 2020) and appeared as early as the work of Zermelo (1929) and Ford (1957) for pairwise comparisons. Hunter (2004) shows that Section 3.1 is necessary and sufficient for the upper-compactness of (4), which guarantees the existence of a maximizer in the interior of the probability simplex. In fact, when an interior maximizer exists, it is also unique, since Section 3.1 implies the following weaker condition, which guarantees the strict concavity of (4). {assumption}[Connectivity] In any partition of $[m]$ into two nonempty subsets $S$ and $S^{C}$, some $j\in S$ and and some $k\in S^{C}$ appear in the same choice set $S_{i}$ for some $i$. The intuitions provided by Ford (1957) and Hunter (2004) are helpful for understanding Assumptions 3.1 and 3.1. If items from some $S\subsetneq[m]$ are never compared with those in $S^{C}$, i.e., never appeared together in any choice set $S_{i}$, it is impossible to rank across the two subsets. In this case, we can rescale the relative weights of $S$ and $S^{C}$ of an interior maximizer and obtain another maximizer. On the other hand, if items in $S$ are always preferred to those in $S^{C}$, we can increase the likelihood by scaling $s_{j}$ for items $j\in S^{C}$ towards 00, and no maximizer in the interior of the probability simplex exists. Nevertheless, a boundary solution can still exist. This case turns out to be important in the present work: in the equivalent matrix balancing problem, it corresponds to the slowdown regime of Sinkhorn’s algorithm, where scalings diverge but the scaled matrix converges (Section 5).
Section 3.1 also has a concise graph-theoretic interpretation. Define the weighted undirected comparison graph $G_{c}$ on $m$ vertices with adjacency matrix $A^{c}$ given by
${A}^{c}_{jk}$ =$=\begin{cases}0&j=k\\ |\{i\mid j,k\in S_{i}\}|&j\neq k,\end{cases}$
In other words, there is an undirected edge between $j$ and $k$ if and only if they are both in some choice set, with the edge weight equal to the number of their co-occurrences. We can verify that Section 3.1 requires $G_{c}$ to be connected.
Under these standard assumptions, previous works have studied the statistical efficiency of the MLE (Hajek et al. 2014, Shah et al. 2015, Seshadri et al. 2020) as well as the computational efficiency of the MM algorithm (Vojnovic et al. 2020). In both cases, the algebraic connectivity of $G_{c}$, quantified by the second smallest eigenvalue of the graph Laplacian of $G_{c}$, plays an important role. See Section 8 for details. In Section 5, we focus on the classic Sinkhorn’s algorithm for matrix balancing, and show that its convergence is also characterized by the algebraic connectivity of a bipartite graph. In Section 10, we extend our framework to settings where Section 3.1 does not hold, by introducing Gamma priors on $s_{j}$, inspired by Caron and Doucet (2012) and Maystre and Grossglauser (2017). These priors regularize the maximum likelihood problem and guarantee that a unique solution in the interior of the probability simplex always exists. They also speed up the associated iterative algorithm by improving the algebraic connectivity.
3.2 The Canonical Matrix Balancing Problem
Matrix balancing is a classic problem that shows up in a wide range of disciplines. See Section 11 for a short survey on some applications. The underlying mathematical problem can be stated concisely in matrix form as:
Given positive vectors $p\in\mathbb{R}_{++}^{n},q\in\mathbb{R}_{++}^{m}$with $\sum_{i}p_{i}=\sum_{j}q_{j}$ and non-negative matrix $A\in\mathbb{R}_{+}^{n\times m}$, find positive diagonal matrices $D^{1}$, $D^{0}$ satisfying the conditions $D^{1}AD^{0}\cdot\mathbf{1}_{m}=p$ and $D^{0}A^{T}D^{1}\cdot\mathbf{1}_{n}=q$.
We henceforth refer to the above as the “canonical” matrix balancing problem. It seeks positive row and column scalings of an (entry-wise) non-negative rectangular matrix $A$ such that the scaled matrix has positive target row and column sums $p$ and $q$. Other variants of the problem replace the row and column sums (the 1-norm) with other norms (Bauer 1963, Ruiz 2001). Note that for any $c>0$, $(D^{0}/c,cD^{1})$ is also a solution whenever $(D^{0},D^{1})$ is. A finite positive solution $(D^{0},D^{1})$ to the canonical matrix balancing problem is often called a direct scaling.
The structure of the matrix balancing problem suggests a simple iterative scheme: starting from any initial positive diagonal $D^{0}$, invert $D^{1}AD^{0}\mathbf{1}_{m}=p$ using $p/(AD^{0}\mathbf{1}_{m})$ to update $D^{1}$. Then invert $D^{0}A^{T}D^{1}\mathbf{1}_{n}=q$ using $q/(A^{T}D^{1}\mathbf{1}_{n})$ to compute the new estimate of $D^{0}$, and repeat the procedure. Here divisions involving two vectors of the same length are entry-wise. This simple scheme is precisely Sinkhorn’s algorithm, described in Algorithm 1, where vectors $d^{0},d^{1}$ are the diagonals of $D^{0},D^{1}$.
Algorithm 1 Sinkhorn’s Algorithm
  Input: $A,p,q,\epsilon_{\text{tol}}$.
  initialize $d^{0}\in\mathbb{R}_{++}^{m}$
  repeat
     $d^{1}\leftarrow p/(Ad^{0})$
     $d^{0}\leftarrow q/({A}^{T}d^{1})$
     ϵitalic-$\epsilon\leftarrow$ update of $(d^{0},d^{1})$
  until $\epsilon<\epsilon_{\text{tol}}$
An important dichotomy occurs depending on whether the entries of $A$ are strictly positive. If $A$ contains no zero entries, then direct scalings and a unique scaled matrix $D^{1}AD^{0}$ always exist (Sinkhorn 1964). Moreover, Sinkhorn’s algorithm converges linearly (Franklin and Lorenz 1989). On the other hand, when $A$ contains zero entries, the canonical problem becomes more complicated. Additional conditions are needed to guarantee meaningful solutions, and the convergence behavior of Sinkhorn’s algorithm is less understood. Well-posedness of the matrix balancing problem has been studied by Brualdi (1968), Sinkhorn (1974), Pukelsheim and Simeone (2009), among others, who characterize the following equivalent existence conditions. {assumption}[Strong Existence] (a) There exists a non-negative matrix $A^{\prime}\in\mathbb{R}_{+}^{n\times m}$ with the same zero patterns as $A$ and with row and column sums $p$ and $q$. Or, equivalently,
(b) For every pair of sets of indices $N\subsetneq[n]$ and $M\subsetneq[m]$ such that $A_{ij}=0$ for $i\notin N$ and $j\in M$, $\sum_{i\in N}p_{i}\geq\sum_{j\in M}q_{j}$, with equality iff $A_{ij}=0$ for all $i\in N$ and $j\notin M$ as well. It is well-known in the matrix balancing literature that the above two conditions are equivalent, and that a positive finite solution $(D^{0},D^{1})$ to the canonical problem exists iff they hold. See, for example, Theorem 6 in Pukelsheim and Simeone (2009). Clearly, Algorithm 1(a) is the minimal necessary condition when a solution to the matrix balancing problem exists. Algorithm 1(b) is closely connected to conditions for perfect matchings in bipartite graphs (Hall 1935). In flow networks (Gale et al. 1957, Ford and Fulkerson 1956, 1957), it is a capacity constraint that guarantees the maximum flow on a bipartite graph with source and sink is equal to $\sum_{i}p_{i}=\sum_{j}q_{j}$ and with positive flow on every edge (Idel 2016). The bipartite graph, denoted by $G_{b}$, is related to $A$ through its adjacency matrix $A^{b}\in\mathbb{R}^{(n+m)\times(n+m)}$ given by
$A^{b}:=\begin{bmatrix}\mathbf{0}&{A}\\ {A}^{T}&\mathbf{0}\end{bmatrix}.$
See Section 8 for a detailed discussion. The structure of $G_{b}$ turns out to be important for the linear convergence rate of Sinkhorn’s algorithm (Section 5).
On the other hand, the necessary and sufficient condition for uniqueness of finite scalings essentially requires that $A$ is not block-diagonal, and guarantees that $G_{b}$ is connected. {assumption}[Uniqueness] $D^{0}$ and $D^{1}$ are unique modulo normalization iff $A$ is indecomposable, i.e., there does not exist permutation matrices $P,Q$ such that $PAQ$ is block diagonal.
With a proper introduction to both problems, we are now ready to formally establish the equivalence between Luce choice model estimation and matrix balancing in the next section. In Section 5, we return to Sinkhorn’s algorithm for the matrix balancing problem, and settle important open problems concerning its linear convergence for non-negative $A$. In Section 9, we discuss further connections of matrix balancing and Sinkhorn’s algorithm to choice modeling and optimization.
4 Connecting Choice Modeling and Matrix Balancing
In this section, we formally establish the connection between choice modeling and matrix balancing. We show that maximizing the log-likelihood (4) is equivalent to solving a canonical matrix balancing problem. We also precisely describe the correspondence between the relevant conditions in the two problems. In view of this equivalence, we show that Sinkhorn’s algorithm, when applied to estimate Luce choice models, is in fact a parallelized generalization of the classic iterative algorithm for choice models, dating back to Zermelo (1929), Dykstra (1956), Ford (1957), and studied extensively also by Hunter (2004), Vojnovic et al. (2020).
4.1 Maximum Likelihood Estimation of Luce Choice Model as Matrix Balancing
The optimality conditions for maximizing the log-likelihood (4) for each $s_{j}$ are given by
$\partial_{s_{j}}\ell(s)=\sum_{j\mid(j,S_{i})}\frac{1}{s_{j}}-\sum_{i\mid j\in S_{i}}\frac{1}{\sum_{k\in S_{i}}s_{k}}$ $=0.$
Multiplying by $s_{j}$ and dividing by $1/n$, we have
$\frac{W_{j}}{n}$ $=\frac{1}{n}\sum_{i\mid j\in S_{i}}\frac{s_{j}}{\sum_{k\in S_{i}}s_{k}},$ (5)
where $W_{j}:=|\{i\mid(j,S_{i})\}|$ is the number of observations where $j$ is selected.
Note that in the special case where $S_{i}\equiv[n]$, i.e., every choice set contains all items, the MLE simply reduces to the familiar empirical frequencies $\hat{s}_{j}={W_{j}}/{n}$. However, when the choice sets $S_{i}$ vary, no closed form solution to (5) exists, which is the primary motivation behind the long line of works on the algorithmic problem of solving (5).
With varying $S_{i}$, we can interpret the optimality condition as requiring the observed frequency of $j$ being chosen (left hand side) be equal to the conditional expected probability of $j$ being chosen among all observations $i$ where it is part of the choice set $S_{i}$ (right hand side). In addition, note that since the optimality condition in (5) only involves the frequency of selection, distinct datasets could yield the same optimality condition and hence the same MLEs. For example, suppose that two alternatives $j$ and $k$ both appear in choice sets $S_{i}$ and $S_{i^{\prime}}$, with $j$ selected in $S_{i}$ and $k$ selected in $S_{i^{\prime}}$. Then switching the choices in $S_{i}$ and $S_{i^{\prime}}$ does not alter the likelihood and optimality conditions. This feature holds more generally with longer cycles of items and choice sets, and can be viewed as a consequence of the context-independent nature of Luce’s choice axiom. In some sense, it is also the underpinning of many works in economics that estimate choice models based on marginal sufficient statistics. A famous example is Berry et al. (1995), which estimates consumer preferences using only aggregate market shares of products.
In practice, the choice sets $S_{i}$ of many observations may be identical to each other. Because (5) only depends on the total winning counts of items, we may aggregate over observations with the same $S_{i}$:
$\sum_{i\mid j\in S_{i}}\frac{s_{j}}{\sum_{k\in S_{i}}s_{k}}$ $=\sum_{i^{\prime}\mid j\in S^{\ast}_{i^{\prime}}}R_{i^{\prime}}\cdot\frac{s_{j}}{\sum_{k\in S^{\ast}_{i^{\prime}}}s_{k}},$
where each $S_{i^{\prime}}^{\ast}$ is a unique choice set that appears in $R_{i^{\prime}}\geq 1$ observations, for $i^{\prime}=1,\dots,n^{\ast}\leq n$. By construction, $\sum_{i^{\prime}=1}^{n^{\ast}}R_{i^{\prime}}=n$. Note, however, that the selected item could vary across different appearances of $S_{i}^{\ast}$, but the optimality conditions only involve each item’s winning count $W_{j}$. From now on, we assume this reduction and drop the ${}^{\ast}$ superscript. In other words, we assume that we observe $n$ unique choice sets, and choice set $S_{i}$ has multiplicity $R_{i}$. The resulting problem has optimality conditions
${W_{j}}$ $=\sum_{i\mid j\in S_{i}}{R_{i}}\cdot\frac{s_{j}}{\sum_{k\in S_{i}}s_{k}}.$ (6)
We are now ready to reformulate (6) as a canonical matrix balancing problem. Define $p\in\mathbb{R}^{n}$ as $p_{i}=R_{i}$, i.e., the number of times choice set $S_{i}$ appears in the data. Define $q\in\mathbb{R}^{m}$ as $q_{j}={W_{j}}$, i.e., the number of times item $j$ was selected in the data. By construction we have $\sum_{i}p_{i}=\sum_{j}q_{j}$, and $p_{i},q_{j}>0$ whenever Section 3.1 holds.
Now define the $n\times m$ binary matrix $A$ by $A_{ij}=1\{j\in S_{i}\}$, so the $i$-th row of $A$ is the indicator of which items appear in the (unique) choice set $S_{i}$, and the $j$-th column of $A$ is the indicator of which choice sets item $j$ appears in. We refer to this $A$ constructed from a choice dataset as the participation matrix. By construction, $A$ has distinct rows, but may still have identical columns. If necessary, we can also remove repeated columns by “merging” items and their win counts. Their estimated scores can be computed from the score of the merged item proportional to their respective win counts.
Let $D^{0}\in\mathbb{R}^{m\times m}$ be the diagonal matrix with $D_{j}^{0}=s_{j}$ and $D^{1}\in\mathbb{R}^{n\times n}$ be the diagonal matrix with $D_{i}^{1}={R_{i}}/{\sum_{k\in S_{i}}s_{k}}$, and define the scaled matrix
$\hat{A}$ $:=D^{1}AD^{0}.$ (7)
The matrices $D^{1}$ and $D^{0}$ are scalings of rows and columns of $A$, respectively, and
$\hat{A}_{ij}=\frac{R_{i}}{\sum_{k\in S_{i}}s_{k}}\cdot 1\{j\in S_{i}\}\cdot s_{j}.$
The key observation is that the optimality condition (6) can be rewritten as
$\hat{A}^{T}\mathbf{1}_{n}$ $=q.$ (8)
Moreover, by construction $\hat{A}$ also satisfies
$\hat{A}\mathbf{1}_{m}$ $=p.$ (9)
Therefore, if $s_{j}$’s satisfy the optimality conditions for maximizing (4), then $D^{0},D^{1}$ defined above solve the matrix balancing problem in Equations 7, 8 and 9. Moreover, the converse is also true, and we thus establish the equivalence between choice maximum likelihood estimation and matrix balancing. All omitted proofs appear in Section 13.
Theorem 4.1
Let $p\in\mathbb{R}^{n}$ with $p_{i}=R_{i}$, $q\in\mathbb{R}^{m}$ with $q_{j}=W_{j}$, and $A\in\mathbb{R}^{n\times m}$ with $A_{ij}=1\{j\in S_{i}\}$ be constructed from the choice dataset. Then $D^{0},D^{1}>0$ with $\sum_{j}D_{j}^{0}=1$ solves the matrix balancing problem
$\begin{split}\hat{A}&=D^{1}AD^{0}\\ \hat{A}\mathbf{1}_{m}&=p\\ \hat{A}^{T}\mathbf{1}_{n}&=q\end{split}$ (10)
if and only if $s\in\Delta_{m}$ with $s_{j}=D^{0}_{j}$ satisfies the optimality condition (6) of the ML estimation problem.
In particular, (4) has a unique maximizer $s$ in the interior of the probability simplex if and only if (10) has a unique normalized solution $D^{0}$ as well. The next question, naturally, is then how Section 3.1 and Section 3.1 for choice modeling are connected to Algorithm 1 and Algorithm 1 for matrix balancing.
Theorem 4.2
Let (A,p,q) be constructed from the choice dataset as in Theorem 4.1, with $p,q>0$. Section 3.1 is equivalent to Algorithm 1. Furthermore, Section 3.1 holds if and only if $(A,p,q)$ satisfy Algorithm 1 and $A$ satisfies Algorithm 1.
Thus when the ML estimation problem is cast as a matrix balancing problem, Algorithm 1 exactly characterizes the gap between Section 3.1 and Section 3.1. We provide some intuition for Theorem 4.2. When we construct a triplet $(A,p,q)$ from a choice dataset, with $p$ the numbers of appearances of unique choice sets and $q$ the winning counts, Algorithm 1 precludes the possibility of partitioning the items into two subsets that never get compared with each other, i.e., Section 3.1. Then Algorithm 1 requires that whenever a strict subset $M\subsetneq[m]$ of objects only appear in a strict subset $N\subsetneq[n]$ of the observations, their total winning counts are strictly smaller than the total number of these observations, i.e., there is some object $j\notin M$ that is selected in $S_{i}$ for some $i\in N$, which is required by Section 3.1.
Interestingly, while Section 3.1 requires the directed comparison graph, defined by the $m\times m$ matrix of counts of item $j$ being chosen over item $k$, to be strongly connected, the corresponding conditions for the equivalent matrix balancing problem concern the $n\times m$ participation matrix $A$ and positive vectors $p,q$, which do not explicitly encode the specific choice of each observation. This apparent discrepancy is due to the fact that $(A,p,q)$ form the sufficient statistics of the Luce choice model. In other words, there can be more than one choice dataset with the same optimality condition (6) and $(A,p,q)$ defining the equivalent matrix balancing problem.
Remark. This feature where “marginal” quantities constitute the sufficient statistics of a parametric model is an important one that underlies many works in economics and statistics (Kullback 1997, Stone 1962, Good 1963, Birch 1963, Theil 1967, Fienberg 1970, Berry et al. 1995, Fofana et al. 2002, Maystre and Grossglauser 2017). It makes the task of estimating a joint model from marginal quantities feasible. This feature is useful because in many applications, only marginal data is available due to high sampling cost or privacy reasons.
Having formulated a particular matrix balancing problem from the estimation problem given choice data, we may ask how one can go in the other direction. In other words, when/how can we construct a “choice dataset” whose sufficient statistics is a given triplet $(A,p,q)$? First off, for $(A,p,q)$ to be valid sufficient statistics of a Luce choice model, $p,q$ need to be positive integers. Moreover, $A$ has to be a binary matrix, with each row containing at least two non-zero elements (valid choice sets have at least two items). Given such a $(A,p,q)$ satisfying Assumptions 1 and 1, a choice dataset can be constructed efficiently. Such a procedure is described, for example, in Kumar et al. (2015), where $A$ is motivated by random walks on a graph instead of matrix balancing (Section 9). Their construction relies on finding the max flow on the bipartite graph $G_{b}$. For rational $p,q$, this maximum flow can be found efficiently in polynomial time (Balakrishnan et al. 2004, Idel 2016). Moreover, the maximum flow implies a matrix $A^{\prime}$ satisfying Algorithm 1(a), thus providing a feasibility certificate for the matrix balancing problem as well.
We have thus closed the loop and fully established the equivalence of the maximum likelihood estimation of Luce choice models and the canonical matrix balancing problem.
Corollary 4.3
There is a one-to-one correspondence between classes of maximum likelihood estimation problems with the same optimality conditions (6) and canonical matrix balancing problems with $(A,p,q)$, where $A$ is a valid participation matrix and $p,q>0$ have integer entries.
We next turn our attention to the algorithmic connections between choice modeling and matrix balancing.
4.2 Algorithmic Connections between Matrix Balancing and Choice Modeling
Given the equivalence between matrix balancing and choice modeling, we can naturally consider applying Sinkhorn’s algorithm to maximize (4). In this case, one can verify that the updates in each full iteration of Algorithm 1 reduce algebraically to
$s_{j}^{(t+1)}$ $=W_{j}/\sum_{i\mid j\in S_{i}}\frac{R_{i}}{\sum_{k\in S_{i}}s_{k}^{(t)}}$ (11)
in the $t$-th iteration. Comparing (11) to the optimality condition in (6), which we recall is given by
$W_{j}$ $=\sum_{i\mid j\in S_{i}}R_{i}\frac{s_{j}}{\sum_{k\in S_{i}}s_{k}}=s_{j}\cdot\sum_{i\mid j\in S_{i}}\frac{R_{i}}{\sum_{k\in S_{i}}s_{k}},$
we can therefore interpret the iterations as simply dividing the winning count $W_{j}$ by the coefficient of $s_{j}$ on the right repeatedly, in the hope of converging to a fixed point. A similar intuition was given by Ford (1957) in the special case of pairwise comparisons. Indeed, the algorithm proposed by Ford (1957) is a cyclic variant of (11) applied to pairwise comparisons. However, this connection is mainly algebraic, as the optimality condition in Ford (1957) does not admit a reformulation as the matrix balancing problem in (10).
In Section 9, we provide further discussions on the connections of Sinkhorn’s algorithm to existing frameworks and algorithms in the choice modeling literature, and connect it to distributed optimization as well. We demonstrate that many existing algorithms for Luce choice model estimation are in fact special cases or analogs of Sinkhorn’s algorithm. These connections also illustrate the many interpretations of Sinkhorn’s algorithm, e.g., as a distributed optimization algorithm as well as a minorization-majorization (MM) algorithm. However, compared to most algorithms for choice modeling discussed in this work, Sinkhorn’s algorithm is more general as it applies to non-binary $A$ and non-integer $p,q$, and has the critical advantage of being paralellized and distributed, hence more efficient in practice.
The mathematical and algorithmic connections between matrix balancing and choice modeling we establish in this paper allow the transfusion of ideas in both directions. For example, inspired by regularized maximum likelihood estimation (Maystre and Grossglauser 2017), we propose a regularized version of Sinkhorn’s algorithm in Section 10, which is guaranteed to converge even when the original Sinkhorn’s algorithm does not converge. Moreover, the importance of algebraic connectivity in quantifying estimation and computation efficiency in choice modeling motivates us to solve some important open problems on the convergence of Sinkhorn’s algorithm. We turn to this topic next.
5 Linear Convergence of Sinkhorn’s Algorithm for Non-negative Matrices
In this section, we turn our focus to matrix balancing and study the global and asymptotic linear convergence rates of Sinkhorn’s algorithm for general non-negative matrices and positive marginals. We first discuss relevant quantities and important concepts before presenting the convergence results in Section 5.2 and Section 5.3. Throughout, we use superscript $(t)$ to denote quantities after $t$ iterations of Sinkhorn’s algorithm.
5.1 Preliminaries
Table 1: Summary of some convergence results on Sinkhorn’s algorithm. In Franklin and Lorenz (1989), $\kappa(A)=\frac{\theta(A)^{1/2}-1}{\theta(A)^{1/2}+1}$, where $\theta(A)$ is the diameter of $A$ in the Hilbert metric. The norm in Knight (2008) is not explicitly specified, and $\sigma_{2}(\hat{A})$ denotes the second largest singular value of the scaled doubly stochastic matrix $\hat{A}$. The bound in Altschuler et al. (2017) was originally stated as $\|r^{(t)}-p\|_{1}\leq\epsilon^{\prime}$ in $t=O(\epsilon^{\prime-2}\log(\frac{\sum_{ij}A_{ij}}{\min_{ij}A_{ij}}))$ iterations. The result in Léger (2021) applies more generally to couplings of probability distributions. In view of Pinsker’s inequality, it implies the bound in Altschuler et al. (2017) but with a constant that is finite even when $A$ has zero entries. In the bound in Knight (2008) and our asymptotic result, the $\lambda+\epsilon$ denotes an asymptotic rate, with the bound valid for any $\epsilon>0$ and all $t$ sufficiently large. In our global bound, the linear rate $\lambda_{-2}(\mathcal{L})$ is the second smallest eigenvalue of the Laplacian of the bipartite graph defined by $A$ (see Section 8), $l=\min\{\max_{j}(A^{T}\mathbf{1}_{n})_{j},\max_{i}(A\mathbf{1}_{m})_{i}\}$, $c_{B}=\exp(-4B)$, and $B$ is a bound on the initial sub-level set, which is finite if and only if Algorithm 1 holds.
convergence statement $\lambda$ $A$ $p,q$
Franklin and Lorenz (1989) $d_{\text{Hilbert}}(r^{(t)},p)\leq\lambda^{t}d_{\text{Hilbert}}(r^{(0)},p)$ $\kappa^{2}(A)$ $A>0$, rectangular uniform
Luo and Tseng (1992) $g(u^{(t)},v^{(t)})-g^{\ast}\leq\lambda^{t}(g(u^{(0)},v^{(0)})-g^{\ast})$ unknown $A\geq 0$, rectangular general
Knight (2008) $\|D_{t+1}^{0}-D^{0}\|_{\ast}\leq(\lambda+\epsilon)\|D_{t}^{0}-D^{0}\|_{\ast}$ $\sigma_{2}^{2}(\hat{A})$ $A\geq 0$, square uniform
Pukelsheim and Simeone (2009) $\|r^{(t)}-p\|_{1}\rightarrow 0$ no rate $A\geq 0$, rectangular general
Altschuler et al. (2017) $\|r^{(t)}-p\|_{1}\leq c\sqrt{\frac{\lambda}{t}}$ $\log(\frac{\sum_{ij}A_{ij}}{\min_{ij}A_{ij}})$ $A>0$, rectangular general
Léger (2021) $D_{\text{KL}}(r^{(t)}\|p)\leq\frac{\lambda}{t}$ $D_{\text{KL}}(\hat{A}\|A)$ $A\geq 0$, continuous general
current work, asymptotic $\|\frac{r^{(t+1)}}{\sqrt{p}}-\sqrt{p}\|_{2}\leq(\lambda+\epsilon)\|\frac{r^{(t)}}{\sqrt{p}}-\sqrt{p}\|_{2}$ $\lambda_{2}(\tilde{A}^{T}\tilde{A})$ $A\geq 0$, rectangular general
current work, global $g(u^{(t)},v^{(t)})-g^{\ast}\leq\lambda^{t}(g(u^{(0)},v^{(0)})-g^{\ast})$ $1-c_{B}\lambda_{-2}(\mathcal{L})/l$ $A\geq 0$, rectangular general
We start with the optimization principles associated with matrix balancing and Sinkhorn’s algorithm. Consider the following KL divergence (relative entropy) minimization problem
$\begin{split}\min_{\hat{A}\in\mathbb{R}^{n\times m}_{+}}&D_{\text{KL}}(\hat{A}\|A)\\ \hat{A}\mathbf{1}_{m}&=p\\ \hat{A}^{T}\mathbf{1}_{n}&=q.\end{split}$ (12)
It is well-known that solutions $\hat{A}=D^{1}AD^{0}$ to the matrix balancing problem with $(A,p,q)$ are minimizers of (12) (Ireland and Kullback 1968, Bregman 1967a). Moreover, Sinkhorn’s algorithm can be interpreted as a block coordinate descent type algorithm applied to minimize the following dual problem of (12):
$g(d^{0},d^{1}):=(d^{1})^{T}Ad^{0}-\sum_{i=1}^{n}p_{i}\log d^{1}_{i}-\sum_{j=1}^{m}q_{j}\log d^{0}_{j},$ (13)
Luo and Tseng (1992) study the linear convergence of block coordinate descent algorithms. Their result implies that the convergence of Sinkhorn’s algorithm, measured in terms of the optimality gap of $g$, is linear with some implicit rate $\lambda>0$, as long as finite positive scalings $D^{0},D^{1}$ exist for the matrix balancing problem. Minimizers $d^{0},d^{1}$ of (13) precisely give the diagonals of $D^{0},D^{1}$. The function $g$, known to be a potential function of Sinkhorn’s algorithm, also turns out to be crucial in quantifying the global linear convergence rate in the present work.
Remark. Interestingly, minimizing (13) is in fact equivalent to maximizing the log-likelihood function $\ell(s)$ in (4) for valid $(A,p,q)$, because $\min_{d^{1}}g(d^{0},d^{1})=-\ell(d^{0})+c$ for some $c>0$. Moreover, the optimality condition of minimizing $g$ with respect to $d^{0}$ reduces to the optimality condition (6). A detailed discussion can be found in Section 9.4. This connection relates choice modeling and matrix balancing from an optimization perspective.
Although convergence results on Sinkhorn’s algorithm are abundant, the recent work of Léger (2021) stands out as the first explicit global convergence result applicable to general non-negative matrices, with a sub-linear $\mathcal{O}(1/t)$ bound on the KL divergence with respect to target marginals. It implies the bounds in Chakrabarty and Khanna (2021), Altschuler et al. (2017) but with a constant that is finite even when $A$ has zero entries. The result in Léger (2021) applies more generally to couplings of continuous probability distributions, but when restricted to the discrete matrix balancing problem, it holds under the following equivalent conditions that are weaker than Algorithm 1. {assumption}[Weak Existence]
(a) There exists a non-negative matrix $A^{\prime}\in\mathbb{R}_{+}^{n\times m}$ that inherits all zeros of $A$ and has row and column sums $p$ and $q$. Or, equivalently,
(b) For every pair of sets of indices $N\subsetneq[n]$ and $M\subsetneq[m]$ such that $A_{ij}=0$ for $i\notin N$ and $j\in M$, $\sum_{i\in N}p_{i}\geq\sum_{j\in M}q_{j}$.
The equivalence of these conditions follows from Theorem 4 in Pukelsheim and Simeone (2009), which also shows that they are the minimal requirements for the convergence of Sinkhorn’s algorithm. Table 1(a) precisely guarantees that the optimization problem (12) is feasible and bounded. It relaxes Algorithm 1(a) by allowing additional zeros in the matrix $A^{\prime}$. Similarly, Table 1(b) relaxes Algorithm 1(b) by allowing equality between $\sum_{i\in N}p_{i}$ and $\sum_{j\in M}q_{j}$ even when $M,N$ do not correspond to a block-diagonal structure.
The distinction between Algorithm 1 and Table 1 is crucial for the matrix balancing problem and Sinkhorn’s algorithm. Recall that Algorithm 1 guarantees the matrix balancing problem has a solution $(D^{0},D^{1})$, and $D^{1}AD^{0}$ is always a solution to (12). On the other hand, the weaker condition Table 1 guarantees that (12) has a solution $\hat{A}$. If indeed $\hat{A}$ has additional zeros relative to $A$, then no direct (finite and positive) scaling $(D^{0},D^{1})$ exists such that $\hat{A}=D^{1}AD^{0}$. However, the sequence of scaled matrices $\hat{A}^{(t)}$ from Sinkhorn’s algorithm still converges to $\hat{A}$. In this case, the matrix balancing problem is said to have a limit scaling, where some entries of $d^{0},d^{1}$ in Sinkhorn iterations approach 0 or $\infty$, resulting in additional zeros in $\hat{A}$. Below we give an example adapted from Pukelsheim and Simeone (2009), where $p,q=(3,3)$ and the scaled matrices $\hat{A}^{(t)}$ converge but no direct scaling exists:
${D^{1}}^{(t)}\begin{bmatrix}3&1\\ 0&2\end{bmatrix}{D^{0}}^{(t)}=\begin{bmatrix}1&0\\ 0&\frac{3t}{2}\end{bmatrix}\begin{bmatrix}3&1\\ 0&2\end{bmatrix}\begin{bmatrix}1&0\\ 0&\frac{1}{t}\end{bmatrix}\rightarrow\begin{bmatrix}3&0\\ 0&3\end{bmatrix}.$
Given these discussions, it is therefore important to clarify the convergence behaviors of Sinkhorn’s algorithm in different regimes. In particular, it remains to reconcile the gap between the implicit linear convergence result of Luo and Tseng (1992) under strong existence, and the quantitative sub-linear bound of Léger (2021) under weak existence. Furthermore, it remains to provide explicit characterizations of both the global and asymptotic (local) rates when Sinkhorn’s algorithm does converge linearly.
Our results in this section provide answers to these questions. We show that the $\mathcal{O}(1/t)$ rate can be sharpened to a global $\mathcal{O}(\lambda^{t})$ bound if and only if the weak existence condition (Table 1) is replaced by the strong existence condition (Algorithm 1). Moreover, we provide an explicit global linear convergence rate $\lambda$ in terms of the algebraic connectivity, revealing the structure-dependent nature of Sinkhorn’s algorithm for problems with non-negative matrices. This generalizes the implicit result of Luo and Tseng (1992) and sheds light on how different assumptions impact Sinkhorn’s convergence, which is explicitly reflected in the constants of the bound. Going further, we characterize the sharp asymptotic rate of linear convergence in terms of the second largest singular value of $\mathcal{D}(1/\sqrt{p})\cdot\hat{A}\cdot\mathcal{D}(1/\sqrt{q})$, where $\mathcal{D}$ denotes the diagonalization of a vector. This asymptotic rate reduces to that given by Knight (2008) for $m=n$ and uniform $p,q$.
The choice of convergence measure is important, and previous works have used different convergence measures. First note that after each iteration in Algorithm 1, the column constraint is always satisfied: ${A}^{(t)}\mathbf{1}_{n}=q$, where ${A}^{(t)}$ is the scaled matrix after $t$ iterations. Léger (2021) uses the KL divergence $D_{\text{KL}}(r^{(t)}\|p)$ between the row sum $r^{(t)}={A}^{(t)}\mathbf{1}_{m}$ and the target row sum $p$ to measure convergence. Franklin and Lorenz (1989) use the Hilbert projective metric between $r^{(t)}$ and $p$. Pukelsheim and Simeone (2009) and Altschuler et al. (2017) use the $\ell^{1}$ distance, which is upper bounded by the KL divergence via Pinsker’s inequality. Knight (2008) focuses on the convergence of the scaling diagonal matrix $D^{0}=\mathcal{D}(d^{0})$ to the optimal solution line, but does not explicitly specify the norm. Some bounds are a priori and hold globally for all iterations, while others hold locally in a neighborhood of the optimum. We summarize the relevant convergence results in Table 1. Here $\lambda_{-2}(S)$ denotes the second smallest eigenvalue of a real symmetric matrix $S$, and $\lambda_{2}(S)$ the second largest eigenvalue. In our work, we characterize the global linear convergence through the optimality gap of (13), which naturally leads to a bound on $\|r^{(t)}-p\|_{1}$. For the sharp asymptotic rate, we choose to use the $\ell^{2}$ distance $\|r^{(t)}/\sqrt{p}-\sqrt{p}\|_{2}$ in order to exploit an intrinsic orthogonality structure afforded by Sinkhorn’s algorithm. This approach results in a novel analysis compared to Knight (2008) that most explicitly reveals the importance of spectral properties in the rate of convergence.
5.2 Global Linear Convergence
We first present the global linear convergence results. Our analysis starts with the following change of variables to transform the potential function (13):
$u:=\log d^{0},\quad v:=-\log d^{1}.$ (14)
This results in the potential function $g(u,v)$ defined as
$g(u,v):=\sum_{ij}A_{ij}e^{-v_{i}+u_{j}}+\sum_{i=1}^{n}p_{i}v_{i}-\sum_{j=1}^{m}q_{j}u_{j},$ (15)
and we can verify that Sinkhorn’s algorithm is equivalent to the alternating minimization algorithm (Bertsekas 1997, Beck and Tetruashvili 2013) for (15), which alternates between minimizing with respect to $u$ and $v$, holding the other block fixed:
$u_{j}^{(t)}\leftarrow\log\frac{q_{j}}{\sum_{i}A_{ij}e^{-v^{(t-1)}_{i}}},\quad v_{i}^{(t)}\leftarrow\log\frac{p_{i}}{\sum_{j}A_{ij}e^{u^{(t)}_{j}}}.$ (16)
The Hessian $\nabla^{2}g(u,v)$ always has $\mathbf{1}_{m+n}$ in its null space. On the surface, standard linear convergence results for first-order methods, which require strong convexity (or related properties like the Polyak–Lojasiewicz condition) of the objective function, do not apply to $g(u,v)$. However, we show that under strong existence and uniqueness conditions for the matrix balancing problem, $g(u,v)$ is in fact strongly convex when restricted to the subspace
$\mathbf{1}_{m+n}^{\perp}:=\{u\in\mathbb{R}^{m},v\in\mathbb{R}^{n}:(u,v)^{T}\mathbf{1}_{m+n}=0\}.$
As a result, Sinkhorn’s algorithm converges linearly with a rate that depends on the (restricted) condition number of its Hessian.
Before proceeding, we introduce a slew of useful definitions. Let Sinkhorn’s algorithm initialize with $u^{(0)}$, and $v^{(0)}$ given by (16). Define the constant $B$ as
$B:$ $=\sup_{(u,v)}\|(u,v)\|_{\infty}$
subject to ($\text{subject to }(u,v)^{T}$ $\mathbf{1}_{m+n}=0,$
$g(u,v)$ $\leq g(u^{(0)},v^{(0)}).$
In other words, $B$ is the diameter of the initial normalized sub-level set. We will show that $B$ is finite and that it bounds normalized Sinkhorn iterates by the coercivity of $g(u,v)$ under Algorithm 1. We similarly define the normalized optimal solution pair
$(u^{\ast},v^{\ast}):=\arg\min_{(u,v)\in\mathbf{1}_{m+n}^{\perp}}g(u,v),$ (17)
and $g^{\ast}:=g(u^{\ast},v^{\ast})$. Finally, define
$l_{0}:=\max_{j}(A^{T}\mathbf{1}_{n})_{j},\quad l_{1}:=\max_{i}(A\mathbf{1}_{m})_{i},$
which are the Lipschitz constants of the two sub-blocks of $g(u,v)$. Next, define the Laplcian matrix $\mathcal{L}$ of the bipartite graph $G_{b}$ as
$\mathcal{L}:$ $=\begin{bmatrix}\mathcal{D}({A}\mathbf{1}_{m})&-{A}\\ -{A}^{T}&\mathcal{D}({A}^{T}\mathbf{1}_{n})\end{bmatrix},$
and refer to the second smallest eigenvalue $\lambda_{-2}(\mathcal{L})$ as the Fiedler eigenvalue. For details on the graph Laplacian and the Fielder eigenvalue, see Section 8.
Using the above notation, we can now state one of our main contributions to the study of Sinkhorn’s algorithm.
Theorem 5.1 (Global Linear Convergence)
Suppose Algorithm 1 and Algorithm 1 hold. For all $t>0$,
$g(u^{(t+1)},v^{(t+1)})-g^{\ast}\leq(1-c_{B}\frac{\lambda_{-2}(\mathcal{L})}{l})\left(g(u^{(t)},v^{(t)})-g^{\ast}\right),$ (18)
where $c_{B}=e^{-4B}$ and $l=\min\{l_{0},l_{1}\}$.
As a consequence, we have the following bound:
$\|r^{(t)}-p\|_{1}$ $\leq c^{\prime}_{B}e^{-c_{B}\frac{\lambda_{-2}(\mathcal{L})}{\min\{l_{0},l_{1}\}}\cdot t},$
where $(c^{\prime}_{B})^{2}=(8B\sum_{i}p_{i})$.
Theorem 5.1 immediately implies the following iteration complexity bound.
Corollary 5.2 (Iteration Complexity)
Under Algorithm 1 and Algorithm 1, $\|r^{(t)}-p\|_{1}\leq\epsilon$ after
$\mathcal{O}\left(\frac{\min\{l_{0},l_{1}\}}{\lambda_{-2}(\mathcal{L})}\cdot\log(1/\epsilon)\right)$
iterations of Sinkhorn’s algorithm.
Remark. The ability of Sinkhorn’s algorithm to exploit the strong convexity of $g(u,v)$ on $\mathbf{1}_{m+n}^{\perp}$ relies critically on the invariance of $g(u,v)$ under normalization, which is an intrinsic feature of the problem that has been largely set aside in the convergence analysis so far. Recall that $u=\log d^{0}$ and $v=-\log d^{1}$, where $d^{0},d^{1}$ are the diagonals of the scaling $(D^{0},D^{1})$. Scalings are only determined up to multiplication by $(1/c,c)$ for $c>0$, and the translation $(u,v)\rightarrow(u-\log c,v-\log c)$ does not alter the objective value in (15). We may therefore impose an auxiliary normalization $(u,v)^{T}\mathbf{1}_{m+n}=0$, or equivalently $\prod_{j}d^{0}_{j}=\prod_{i}d^{1}_{i}$. This normalization is easily achieved by requiring that after every update of Sinkhorn’s algorithm, a normalization $(d^{0}/c,cd^{1})$ is performed using the normalizing constant
$c=\sqrt{\prod_{j}d^{0}_{j}/\prod_{i}d^{1}_{i}}.$ (19)
See Algorithm 2. Note, however, that this normalization is only a supplementary construction in our analysis. The final convergence result applies to the original Sinkhorn’s algorithm without normalization, since it does not alter the objective value. Normalization of Sinkhorn’s algorithm is discussed in Carlier et al. (2023), although they use the asymmetric condition $u_{0}=0$, which does not guarantee that normalized Sinkhorn iterates stay in $\mathbf{1}_{m+n}^{\perp}$.
Algorithm 2 Normalized Sinkhorn’s Algorithm
  Input: $A,p,q,\epsilon_{\text{tol}}$.
  initialize $d^{0}\in\mathbb{R}_{++}^{m}$
  repeat
     $d^{1}\leftarrow p/(Ad^{0})$
     normalization $(d^{0},d^{1})\leftarrow(d^{0}/c,cd^{1}),c>0$
     $d^{0}\leftarrow q/({A}^{T}d^{1})$
     normalization $(d^{0},d^{1})\leftarrow(d^{0}/c,cd^{1}),c>0$
     ϵitalic-$\epsilon\leftarrow$ update of $(d^{0},d^{1})$
  until $\epsilon<\epsilon_{\text{tol}}$
The proof of Theorem 5.1 then relies crucially on the observation that the Hessian of $g(u,v)$ at $(0,0)$ is precisely the Laplacian $\mathcal{L}$ of the bipartite graph $G_{b}$. Therefore, as $(u,v)$ are bounded throughout the iterations thanks to the coercivity of $g$, the Fiedler eigenvalue of $\mathcal{L}$ quantifies the strong convexity on $\mathbf{1}_{m+n}^{\perp}$. The linear convergence then follows from standard results on block coordinate descent and alternating minimization methods for strongly convex and smooth functions (Beck and Tetruashvili 2013). Typically, the leading eigenvalue of the Hessian quantifies the smoothness (Luenberger et al. 1984). This is given by $2\max\{l_{0},l_{1}\}$ for $\mathcal{L}$. However, for alternating minimization methods, the better smoothness constant $\min\{l_{0},l_{1}\}$ is available. Thus the quantity $\min\{l_{0},l_{1}\}/\lambda_{-2}(\mathcal{L})$ can be interpreted as a type of “condition number” of the graph Laplacian $\mathcal{L}$. When $A$ is positive (not just non-negative), then the strong existence and uniqueness conditions are trivially satisfied, and our results continue to hold with the rate quantified by $\min\{l_{0},l_{1}\}/\lambda_{-2}(\mathcal{L})$.
Remark. The importance of Assumptions 1 and 1 are clearly reflected in the bound (18). First, note that the Fiedler eigenvalue $\lambda_{-2}(\mathcal{L})>0$ iff Algorithm 1 holds (see Section 8). On the other hand, Algorithm 1 guarantees the coercivity of $g$ on $\mathbf{1}_{m+n}^{\perp}$. This property ensures that $B<\infty$, and consequently, that normalized iterates stay bounded by $B$. That Algorithm 1 guarantees $g(u,v)$ is coercive should be compared to the observation by Hunter (2004) that Section 3.1 guarantees the upper compactness (a closely related concept) of the log-likelihood function (4). In contrast, when only the weak existence condition (Table 1) holds, finite minimizer of $g(u,v)$ may not exist, in which case the diameter $B$ of the initial sub-level set may become infinite.
Table 1 corresponds to the “limit scaling” regime of Sinkhorn’s algorithm, where the scaled matrices $A^{(t)}$ are guaranteed to converge to a finite matrix $\hat{A}$ with the desired marginal distributions that solves (12), and may have additional zeros compared to $A$. Under Table 1, Léger (2021) shows the slower $\mathcal{O}(1/t)$ convergence in KL divergence $D_{\text{KL}}(r^{(t)}\|p)$. We now show that this rate is tight, which fully characterizes the following convergence behavior of Sinkhorn’s algorithm: whenever a direct scaling exists for the matrix balancing problem, Sinkhorn’s algorithm converges linearly. If only a limit scaling exists, then convergence deteriorates to $\mathcal{O}(1/t)$.
Theorem 5.3
For general non-negative matrices, Sinkhorn’s algorithm converges linearly iff $(A,p,q)$ satisfy Algorithm 1 and Algorithm 1. The convergence deteriorates to sub-linear iff the weak existence condition Table 1 holds but Algorithm 1 fails.
The regime of sub-linear convergence also has an interpretation in the choice modeling framework. The weak existence condition Table 1, when applied to $(A,p,q)$ constructed from a choice dataset, allows the case where some subset $S$ of items is always preferred over $S^{C}$, which implies, as observed already by the early work of Ford (1957), that the log-likelihood function (4) is only maximized at the boundary of the probability simplex, by shrinking $s_{j}$ for $j\in S^{C}$ towards 0, i.e., $D^{0}_{j}\rightarrow 0$. Incidentally, Bacharach (1965) also refers to the corresponding regime in matrix balancing as “boundary solutions”.
5.3 Sharp Asymptotic Rate
Having established the global convergence of Sinkhorn’s algorithm when finite scalings exist, we now turn to the open problem of characterizing its asymptotic linear convergence rate for non-uniform marginals. Our analysis relies on an intrinsic orthogonality structure of Sinkhorn’s algorithm instead of the auxiliary normalization used to prove the global linear convergence above. Note that unlike the global rate, which depends on $A$, the asymptotic rate now depends on the associated solution $\hat{A}$ (and $p,q$), as expected.
Theorem 5.4 (Sharp Asymptotic Rate)
Suppose $(A,p,q)$ satisfy Algorithm 1 and Algorithm 1. Let $\hat{A}$ be the unique scaled matrix with marginals $p,q$. Then
$\lim_{t\rightarrow\infty}\frac{\|r^{(t+1)}/\sqrt{p}-\sqrt{p}\|_{2}}{\|r^{(t)}/\sqrt{p}-\sqrt{p}\|_{2}}=\lambda_{\infty},$
where the asymptotic linear rate of convergence $\lambda_{\infty}$ is
$\lambda_{\infty}$ $:=\lambda_{2}(\tilde{A}\tilde{A}^{T})=\lambda_{2}(\tilde{A}^{T}\tilde{A})$
$\tilde{A}$ $:=\mathcal{D}(1/\sqrt{p})\cdot\hat{A}\cdot\mathcal{D}(1/\sqrt{q}),$
where $\lambda_{2}(\cdot)$ denotes the second largest eigenvalue.
Intuitively, the dependence of the linear rate of convergence on the second largest eigenvalue of $\tilde{A}^{T}\tilde{A}$ (and $\tilde{A}\tilde{A}^{T}$) is due to the fact that near the optimum $\sqrt{p}$, $\tilde{A}\tilde{A}^{T}$ (which is the Jacobian at $\sqrt{p}$) approximates the first order change in $r^{(t)}/\sqrt{p}$. Normally, the leading eigenvalue quantifies this change. The unique leading eigenvalue of $\tilde{A}\tilde{A}^{T}$ is equal to 1 with eigenvector $\sqrt{p}$, which does not imply contraction. Fortunately, using the quantity $r^{(t)}/\sqrt{p}$ allows us to exploit the following orthogonality structure:
$(r^{(t)}/\sqrt{p}-\sqrt{p})^{T}\sqrt{p}$ $=\sum_{i}(r_{i}^{(t)}-p_{i})=0$
by virtue of Sinkhorn’s algorithm preserving the quantities $r^{(t)T}\mathbf{1}_{n}$ for all $t$. Thus, the residual $r^{(t)}/\sqrt{p}-\sqrt{p}$ is always orthogonal to $\sqrt{p}$, which is both the leading eigenvector and the fixed point of the iteration. The convergence is then controlled by the second largest eigenvalue of $\tilde{A}\tilde{A}^{T}$. This proof approach echoes that of the global linear convergence result in Theorem 5.1, where we also exploit an orthogonality condition to obtain a meaningful bound. In Theorem 5.1 the bound depends on the second smallest eigenvalue of the Hessian, while in Theorem 5.4 the bound depends on the second largest eigenvalue of the Jacobian.
In the special case of $m=n$ and $p=q=\mathbf{1}$, the asymptotic rate in Theorem 5.4 reduces to that in Knight (2008). Note, however, that the convergence metric is different: we use the $\ell^{2}$ norm $\|r^{(t)}/\sqrt{p}-\sqrt{p}\|_{2}$ while Knight (2008) uses an implicit norm that measures the convergence of $D^{0}$ to the solution line due to scale invariance. Our analysis exploits the orthogonality structure of Sinkhorn’s algorithm and more explicitly reveals the dependence of the convergence rate on the spectral structure of the data.
Our results in this section are relevant in several respects. First, we clarify the gap between the $\mathcal{O}(1/t)$ and $\mathcal{O}(\lambda^{t})$ convergence of Sinkhorn’s algorithm: the slowdown happens if and only if Sinkhorn’s algorithm converges but the canonical matrix balancing problem does not have a finite scaling $(D^{0},D^{1})$. This slowdown has been observed in the literature but not systematically studied. Second, we settle open problems and establish the first quantitative global linear convergence result for Sinkhorn’s algorithm applied to general non-negative matrices. We also characterize the asymptotic linear rate of convergence, generalizing the result of Knight (2008) but with a novel analysis. Third, our analysis reveals the importance of algebraic connectivity for the convergence of Sinkhorn’s algorithm. Although an important quantity in the choice modeling literature, algebraic connectivity has not been previously used to in the analysis of Sinkhorn’s algorithm. The importance of algebraic connectivity for Sinkhorn’s algorithm becomes less surprising once we connect it to the distributed optimization literature in Section 9, where it is well-known that the spectral gap of the gossip matrix, which defines the decentralized communication network, governs the rates of convergence.
6 Conclusion
In this paper, we develop extensive connections between matrix balancing and the estimation of a broad class of choice models. In particular, many algorithms in choice modeling can be viewed as special cases or analogs of Sinkhorn’s algorithm for matrix balancing. These connections benefit both disciplines. For choice modeling, they open the door to tools and insights from a rich research area in optimization and numerical linear algebra, potentially leading to new results on the estimation of choice models. For matrix balancing, the connections inspire us to resolve some long-standing open problems on the linear convergence of Sinkhorn’s algorithm for non-negative matrices, revealing the importance of algebraic connectivity and related spectral properties. Moreover, we propose regularization methods for Sinkhorn’s algorithm inspired by works from choice modeling, in order to address existence and convergence issues for matrix balancing. We believe that the connections we establish in this paper are useful for researchers from both domains and can lead to further interesting results.
7 Acknowledgements
This work is supported in part by NSF CAREER Award #2143176. We are very grateful for insightful comments and suggestions from Serina Chang, Patrick Ding, Wenzhi Gao, Guido Imbens, Süleyman Kerimov, Yongchan Kwon, Han Hong, Frederic Koehler, Pavel Shibayev, Ruoxuan Xiong, and Yinyu Ye.